### v1

# Quantitative BMO–BLO Estimates for the Hardy–Littlewood Maximal Function

Abstract.
In this note, we study a quantitative extension of the John-Nirenberg inequality for the Hardy-Littlewood maximal function of a $\operatorname{BMO}$ function. More precisely, for every nonconstant locally integrable function $f$ such that MfMf is not identically infinite, we prove the inequality
$\left(\frac{1}{w(Q)}\int_{Q}\left(\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}\right)^{p}w(x)\,dx\right)^{\frac{1}{p}}\leq c_{n}\,[w]_{A_{\infty}}\,p$
for every cube $Q$, every $1\leq p<\infty$ and every weight $w\in A_{\infty}$, where $w]_{A_{\infty}}$ denotes the Fujii-Wilson $A_{\infty}$ constant. This result extends the classical boundedness $\|Mf\|_{\operatorname{BLO}}\leq C_{n}\|f\|_{\operatorname{BMO}}$ ([BDS, Bennett]). Furthermore, we show that the class $A_{\infty}$ is both necessary and sufficient for this inequality to hold, providing a new characterization of $A_{\infty}$ in terms of the action of the maximal operator on bounded oscillation spaces.
Key words and phrases:
BMO, John-Nirenberg inequality, Muckenhoupt weights, Maximal function
2020 Mathematics Subject Classification:
Primary 30H35, 42B25; Secondary 42B35
This research is supported by the Basque Government through the BERC 2022-2025 program, by the Ministry of Science and Innovation: Grant PRE2021-099091 funded by BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN / AEI / 10.13039/501100011033 and by ESF+.
1. Introduction and statement of the main results
The theory of functions of bounded mean oscillation ($\operatorname{BMO}$) plays a fundamental role in modern harmonic analysis. The classical John-Nirenberg theorem [JN] states that functions of bounded mean oscillation are locally exponentially integrable. More precisely, if $f\in\operatorname{BMO}$, then there exist dimensional constants $c_{1},c_{2}>0$ such that for every cube $Q\subset{\mathbb{R}}^{n}$,
$|\{x\in Q:\,|f(x)-f_{Q}|>t\}|\leq c_{1}e^{-c_{2}t/\|f\|_{\operatorname{BMO}}}|Q|,\qquad t>0,$
where
$\|f\|_{\operatorname{BMO}}=\sup_{Q}\frac{1}{|Q|}\int_{Q}|f(x)-f_{Q}|\,dx$
with $f_{Q}=\frac{1}{|Q|}\int_{Q}f$, and the supremum is taken over all cubes in ${\mathbb{R}}^{n}$ with sides parallel to the coordinate axes. This inequality establishes that bounded mean oscillation implies a form of exponential control stronger than any $L^{p}$ bound. This property also identified the $\operatorname{BMO}$ space as a natural endpoint substitute for $L^{\infty}$ in the scale of Lebesgue spaces and as a central object in the study of singular integrals, interpolation, and partial differential equations.
The Hardy-Littlewood maximal operator
$Mf(x)=\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}|f(y)|\,dy,$
and the sharp maximal operator introduced by Fefferman and Stein [FS],
$M^{\#}f(x)=\sup_{Q\ni x}\frac{1}{|Q|}\int_{Q}|f(y)-f_{Q}|\,dy,$
are key tools in the study of singular integrals, regularity, weighted inequalities, and self-improvement phenomena. These operators satisfy the pointwise inequality $M^{\#}f(x)\leq 2Mf(x)$, and the Fefferman-Stein inequality shows that both operators are equivalent in $L^{p}$ norms. The exponential integrability in the John-Nirenberg inequality may be viewed as an endpoint version of this relation.
Many authors have studied quantitative versions of the John-Nirenberg inequality, including extensions to weighted settings [MRR, PR, OPRR, JCCP], to metric measure spaces [Stromberg, StrombergTorchinsky, FPW, DGL, Kim, GK], and even to more general measures and capacities [ChenSpector, BCRS]. Indeed, the weighted John-Nirenberg, Poincaré, and Poincaré-Sobolev inequalities can be unified within a single theory through the so-called self-improving phenomena (see [FPW, MP, PR, GLP]).
In [Karagulyan], Karagulyan obtained an extension of the John-Nirenberg inequality, later refined by Canto and Pérez [JCCP]. This extension shows that one can recover the exponential decay of the oscillation of any function once it is normalized by its sharp maximal function. More precisely, the main result of [JCCP] established the inequality
$\left(\frac{1}{w(Q)}\int_{Q}\left(\frac{M_{Q}(f-f_{Q})(x)}{M^{\#}f(x)}\right)^{p}w(x)\,dx\right)^{\frac{1}{p}}\leq C_{n}[w]_{A_{\infty}}p,$ (1.1)
for every cube $Q$, every $1\leq p<\infty$ and every weight $w\in A_{\infty}$ (we also refer to [LLO] for a different proof of (1.1)). Here, $M_{Q}$ is the local dyadic maximal operator, and $w]_{A_{\infty}}$ denotes the Fujii-Wilson $A_{\infty}$ constant, introduced in [HP], defined by
$w]_{A_{\infty}}=\sup_{Q}\frac{1}{w(Q)}\int_{Q}M(w\chi_{Q})(x)\,dx.$
In a classical result, Bennett, DeVore, and Sharpley [BDS] proved that the Hardy-Littlewood maximal operator $M$ maps $\operatorname{BMO}$ into itself. Shortly after, Bennett [Bennett] refined this result by proving that $M$ actually maps $\operatorname{BMO}$ into a smaller space, the space of bounded lower oscillation functions ($\operatorname{BLO}$), satisfying
$\|Mf\|_{\operatorname{BLO}}\leq C_{n}\|f\|_{\operatorname{BMO}},$ (1.2)
for every function $f\in\operatorname{BMO}$ such that MfMf is not identically infinite. Introduced by Coifman and Rochberg in [CR], the space $\operatorname{BLO}$ consists of all locally integrable functions such that
$\|f\|_{\operatorname{BLO}}=\sup_{Q}\frac{1}{|Q|}\int_{Q}\left(f(x)-\operatorname*{ess\,inf}_{Q}f\right)\,dx<\infty.$
From the perspective of weight theory, these spaces admit a natural characterization in terms of logarithms of Muckenhoupt weights, as shown in [CR]. In fact, one can write
$\operatorname{BLO}=\{\alpha\log w:w\in A_{1},\ \alpha\in{\mathbb{R}}\},$
and
$\operatorname{BMO}=\{\alpha\log w:w\in A_{\infty},\ \alpha\in{\mathbb{R}}\}.$
This formulation shows the intrinsic link between bounded oscillation spaces and Muckenhoupt weights, and clearly shows the strict inclusion $\operatorname{BLO}\subsetneq\operatorname{BMO}$.
The present note is devoted to proving a quantitative extension of the mapping property (1.2), analogous to (1.1). In particular, we obtain a quantitative version of this inequality in the spirit of the John-Nirenberg inequality. Concretely, we prove the following result.
Theorem $1.$
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite, and let $w\in A_{\infty}$ be a weight. Then, for any cube $Q$ and $1\leq p<\infty$, we have
$\left(\frac{1}{w(Q)}\int_{Q}\left(\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}\right)^{p}w(x)\,dx\right)^{\frac{1}{p}}\leq c_{n}\,[w]_{A_{\infty}}\,p.$ (1.3)
In particular, if $f\in\operatorname{BMO}$, then for each $1\leq p<\infty$ we have
$\left(\frac{1}{w(Q)}\int_{Q}\left(Mf(x)-\operatorname*{ess\,inf}_{Q}Mf\right)^{p}w(x)\,dx\right)^{\frac{1}{p}}\leq c_{n}\,[w]_{A_{\infty}}\,p\,\|f\|_{\operatorname{BMO}},$ (1.4)
for all cubes $Q$.
Remark 1.2.
The assumption that MfMf is not identically infinite cannot be removed. In [BDS], it is shown that this condition is necessary for the validity of the boundedness of $M$ on $\operatorname{BMO}$, and an explicit counterexample is provided by the function $f(x)=\log|x|$. We also refer to the recent work [BGSKM], which characterizes the class of measures $\mu$ for which $M\mu$ is finite almost everywhere.
The linear dependence on $p$ in the right-hand side of (1.3) is essential, as it yields exponential estimates as a direct corollary.
Corollary 1.3.
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. Let $w\in A_{\infty}$ be a weight.
  • There exist dimensional constants $c_{1},c_{2}>0$ such that
    $w\left(\left\{x\in Q:\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}>t\right\}\right)\leq c_{1}e^{-\frac{c_{2}t}{[w]_{A_{\infty}}}}w(Q),$ (1.5)
    for all $t>0$ and all cubes $Q$. That is,
    $\left\|\frac{Mf-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f}\right\|_{\exp L\left(Q,\frac{w(x)\,dx}{w(Q)}\right)}\leq c_{n}[w]_{A_{\infty}}.$
  • There exist dimensional constants $c_{1},c_{2}>0$ such that for every cube $Q$ and for every $\lambda,\gamma>0$ the following good-lambda inequality holds:
    $w\left(\left\{x\in Q:Mf(x)-\operatorname*{ess\,inf}_{Q}Mf>\lambda,M^{\#}f(x)\leq\gamma\lambda\right\}\right)\leq c_{1}e^{-\frac{c_{2}}{[w]_{A_{\infty}}\gamma}}w(Q).$ (1.6)
Remark 1.4.
We note that inequality (1.6) is an improvement of the classical good-lambda estimate of Journé [J, p. 43], since it provides exponential decay.
The previous inequality (1.4), in the particular case $p=1$, tells us that if $w\in A_{\infty}$ then we have the boundedness
$\|Mf\|_{\operatorname{BLO}_{w}}\leq C_{n}[w]_{A_{\infty}}\|f\|_{\operatorname{BMO}},$
where
$\|f\|_{\operatorname{BLO}_{w}}:=\sup_{Q}\frac{1}{w(Q)}\int_{Q}\left(f(x)-\operatorname*{ess\,inf}_{Q}f\right)w(x)\,dx.$
We can also prove that the converse implication holds: if the above inequality is satisfied for all locally integrable functions $f$ for which MfMf is not identically infinite, then the weight $w$ necessarily belongs to the class $A_{\infty}$. This equivalence yields a new quantitative characterization of the $A_{\infty}$ class in terms of the boundedness of the Hardy-Littlewood maximal operator acting on $\operatorname{BMO}$.
Theorem 1.5.
Let $w$ be a weight. Then there exist some dimensional constants $c_{n},C_{n}$ such that
$c_{n}[w]_{A_{\infty}}\leq\sup_{f}\|Mf\|_{\operatorname{BLO}_{w}}\leq C_{n}[w]_{A_{\infty}},$ (1.7)
where the supremum is taken over all functions $f$ such that $\|f\|_{\operatorname{BMO}}=1$ and MfMf is not identically infinite. As a consequence, we have the following quantitative equivalence of the Fujii-Wilson $A_{\infty}$ constant,
$w]_{A_{\infty}}\simeq_{n}\sup_{f}\sup_{Q}\frac{1}{w(Q)}\int_{Q}\left(Mf(x)-\operatorname*{ess\,inf}_{Q}Mf\right)w(x)\,dx.$
As a direct corollary, we obtain that the class $A_{\infty}$ is both necessary and sufficient for the validity of inequality (1.3).
Corollary 1.6.
Let $w$ be a weight.
  • Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. If $w\in A_{\infty}$, then for any cube $Q$ and $1\leq p<\infty$, we have
    $\left(\frac{1}{w(Q)}\int_{Q}\left(\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}\right)^{p}w(x)\,dx\right)^{\frac{1}{p}}\leq c_{n}[w]_{A_{\infty}}p.$
  • Assume that the estimate
    $\left(\frac{1}{w(Q)}\int_{Q}\left(\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}\right)^{p}w(x)\,dx\right)^{\frac{1}{p}}\leq C\,p$ (1.8)
    holds for every nonconstant locally integrable function $f$ such that MfMf is not identically infinite, every $1\leq p<\infty$, and every cube Q. Then $w\in A_{\infty}$ and $w]_{A_{\infty}}\lesssim_{n}C$.
Remark 1.7.
A similar characterization holds for the main result of [JCCP], namely inequality (1.1). The proof follows exactly the same argument as in the corollary above, replacing the use of Theorem 1.5 with [OPRR, Theorem 1.2].
As usual, $C$ denotes a positive constant, possibly varying from line to line. We write $C_{\alpha,\beta,...}$ to denote a constant depending only on $\alpha,\beta,$....
2. The unweighted case
In this section, we present the proof of the main result, Theorem 1.1, in the unweighted setting. We first recall the following theorem from [JCCP, Theorem 5.1], which extends the exponential-type form of (1.1) to the classical (non-dyadic) maximal function. The proof relies on a result of [Conde], which is used to combine estimates over a finite family of different dyadic lattices and then recover the non-dyadic case.
Theorem 2.1.
Let $Q$ be an arbitrary cube and $f$ a nonconstant locally integrable function. Then for any $\lambda>0$ we have
$\left|\left\{x\in Q:\frac{M\left(\left(f-f_{Q}\right)\chi_{Q}\right)(x)}{M^{\#}f(x)}>\lambda\right\}\right|\leq Ce^{-c\lambda}|Q|,$ (2.1)
where $C,c>0$ are dimensional constants.
Using the layer-cake formula, (2.1), and noting that $\Gamma(p+1)^{1/p}\lesssim p$, we obtain the following inequality,
$\left(\frac{1}{|Q|}\int_{Q}\left(\frac{M((f-f_{Q})\chi_{Q})(x)}{M^{\#}f(x)}\right)^{p}\,dx\right)^{\frac{1}{p}}\leq C_{n}p,$ (2.2)
for each cube $Q$, and each $1\leq p<\infty$ (we refer to the proof of Theorem 1.1 below for more details). With this preliminary estimate established, we are now in a position to prove the unweighted version of Theorem $1.$
Proposition $2.$
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. Then, there exists a dimensional constant $c_{n}>0$ such that for every $1\leq p<\infty$, we have
$\left(\frac{1}{|Q|}\int_{Q}\left(\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}\right)^{p}\,dx\right)^{\frac{1}{p}}\leq c_{n}\,p.$
for every cube $Q$.
In the proof, we decompose MfMf into two terms: the local and nonlocal parts. The local contribution is estimated using (2.2), whereas the nonlocal term is treated following ideas from [Bennett].
Proof.
Fix a cube $Q$ and $1\leq p<\infty$. We decompose the maximal function as
$Mf(x)=$ $M((f-f_{3Q})\chi_{3Q}+f_{3Q}\chi_{3Q}+f\chi_{(3Q)^{c}})(x)$
$\leq$ $M((f-f_{3Q})\chi_{3Q})(x)+Mg(x),$
where $g(x):=f_{3Q}\chi_{3Q}(x)+f(x)\chi_{(3Q)^{c}}(x)$, and we denote by $3Q$ the cube with the same center as $Q$ and side length $3\ell(Q)$. Therefore,
$Mf(x)-\operatorname*{ess\,inf}_{Q}Mf\leq$ $M((f-f_{3Q})\chi_{3Q})(x)+Mg(x)-\operatorname*{ess\,inf}_{Q}Mf$
$\leq$ $M((f-f_{3Q})\chi_{3Q})(x)+\left(Mg(x)-\operatorname*{ess\,inf}_{Q}Mf\right)^{+},$
where $a^{+}=\max(a,0)$. This yields the estimate
$\left(\frac{1}{|Q|}\int_{Q}\left(\frac{Mf-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f}\right)^{p}\,dx\right)^{\frac{1}{p}}\leq$ $\left(\frac{1}{|Q|}\int_{Q}\left(\frac{M((f-f_{3Q})\chi_{3Q})}{M^{\#}f}\right)^{p}\,dx\right)^{\frac{1}{p}}$
$+\left(\frac{1}{|Q|}\int_{Q}\left(\frac{(Mg-\operatorname*{ess\,inf}_{Q}Mf)^{+}}{M^{\#}f}\right)^{p}\,dx\right)^{\frac{1}{p}}$
$=:$ $I_{1}+I_{2}.$
By applying (2.2) over the larger cube $3Q$, we obtain
$I_{1}\leq$ $\left(\frac{1}{|Q|}\int_{3Q}\left(\frac{M((f-f_{3Q})\chi_{3Q})(x)}{M^{\#}f(x)}\right)^{p}\,dx\right)^{\frac{1}{p}}$
$\leq$ $\left(\frac{3^{n}}{|3Q|}\int_{3Q}\left(\frac{M((f-f_{3Q})\chi_{3Q})(x)}{M^{\#}f(x)}\right)^{p}\,dx\right)^{\frac{1}{p}}$
$\leq$ $c_{n}p.$
To estimate the second term, we bound MgMg pointwise in $Q$. Let $x\in Q$ and $P$ be a cube such that $x\in P$. We consider two cases. First, if $P\subseteq 3Q$, then
$\frac{1}{|P|}\int_{P}|g(y)|\,dy=|f_{3Q}|\leq\frac{1}{|3Q|}\int_{3Q}|f(y)|\,dy\leq\operatorname*{ess\,inf}_{z\in Q}Mf(z).$
Otherwise, if $P\cap(3Q)^{c}\neq\emptyset$, we consider the smallest cube $\bar{P}$ that contains both $P$ and $3Q$. Since $x\in Q\cap P$, we have $|\bar{P}|\leq c_{n}|P|$. Hence,
$\frac{1}{|P|}\int_{P}|g(y)|\,dy\leq$ $\frac{1}{|P|}\int_{P}|g(y)-f_{\bar{P}}|\,dy+|f_{\bar{P}}|$
$=:$ $II_{1}+II_{2}.$
We can estimate $II_{2}$ trivially,
$II_{2}\leq\frac{1}{|\bar{P}|}\int_{\bar{P}}|f(y)|\,dy\leq\operatorname*{ess\,inf}_{z\in Q}Mf(z).$
Using the definition of $g$, we have
$II_{1}\leq$ $c_{n}\frac{1}{|\bar{P}|}\int_{\bar{P}}|g(y)-f_{\bar{P}}|\,dy$
$=$ $c_{n}\frac{1}{|\bar{P}|}\int_{\bar{P}}|f_{3Q}\chi_{3Q}(y)+f(y)\chi_{(3Q)^{c}}(y)-f_{\bar{P}}|\,dy$
$=$ $c_{n}\frac{1}{|\bar{P}|}\int_{3Q}|f_{3Q}\chi_{3Q}(y)+f(y)\chi_{(3Q)^{c}}(y)-f_{\bar{P}}|\,dy$
$+c_{n}\frac{1}{|\bar{P}|}\int_{\bar{P}\cap(3Q)^{c}}|f_{3Q}\chi_{3Q}(y)+f(y)\chi_{(3Q)^{c}}(y)-f_{\bar{P}}|\,dy$
$=$ $c_{n}\frac{1}{|\bar{P}|}\int_{3Q}|f_{3Q}-f_{\bar{P}}|\,dy+c_{n}\frac{1}{|\bar{P}|}\int_{\bar{P}\cap(3Q)^{c}}|f(y)-f_{\bar{P}}|\,dy$
$\leq$ $c_{n}\frac{1}{|\bar{P}|}|3Q|\frac{1}{|3Q|}\int_{3Q}|f(y)-f_{\bar{P}}|\,dy+c_{n}\frac{1}{|\bar{P}|}\int_{\bar{P}\cap(3Q)^{c}}|f(y)-f_{\bar{P}}|\,dy$
$=$ $c_{n}\frac{1}{|\bar{P}|}\int_{\bar{P}}|f(y)-f_{\bar{P}}|\,dy$
$\leq$ $c_{n}M^{\#}f(x).$
We have proved
$\frac{1}{|P|}\int_{P}|g(y)|\,dy\leq c_{n}M^{\#}f(x)+\operatorname*{ess\,inf}_{z\in Q}Mf(z),$
for each cube $P$ that contains $x$, and thus
$Mg(x)\leq c_{n}M^{\#}f(x)+\operatorname*{ess\,inf}_{z\in Q}Mf(z),$
for every $x\in Q$. Using this estimate, we can bound $I_{2}$,
$I_{2}\leq$ $\left(\frac{1}{|Q|}\int_{Q}\left(\frac{c_{n}M^{\#}f(x)}{M^{\#}f(x)}\right)^{p}\,dx\right)^{\frac{1}{p}}=c_{n}\leq c_{n}p.$
This concludes the proof. ∎
The linear dependence on $p$ in (2.2) (and in (1.3)) implies exponential-type integrability (see [JCCP, Lemma 2.3]).
Corollary 2.3.
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. Then, there exists a dimensional constant $c_{n}>0$ such that
$\left|\left\{x\in Q:\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}>t\right\}\right|\leq 2e^{-c_{n}t}|Q|,$ (2.3)
for all $t>0$ and all cubes $Q$.
3. Weighted estimates and the characterization of $A_{\infty}$
A weight $w\in A_{\infty}$ satisfies
$\frac{w(E)}{w(Q)}\leq 2\left(\frac{|E|}{|Q|}\right)^{\frac{1}{2^{n+1}[w]_{A_{\infty}}}},$
for every cube $Q$ and every measurable set $E\subseteq Q$ (we refer to [HPR]). Using this property of $A_{\infty}$ weights and (2.3), we can prove Corollary 1.3. We have chosen to follow the strategy of first establishing the unweighted case and then passing to the weighted estimate using the above property. This approach avoids relying on the doubling property of $A_{\infty}$ weights, whose doubling constant is known to depend doubly exponentially on $w]_{A_{\infty}}$ (see [HaPa, p. 945] ). Our approach yields linear dependence on $w]_{A_{\infty}}$ in (1.3).
Proof of Corollary 1.3.
Let $w\in A_{\infty}$ and fix a cube $Q$. Let $t>0$ and denote
$E_{t}=\left\{x\in Q:\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}>t\right\}.$
By the previous property of $A_{\infty}$ weights and (2.3),
$w\left(E_{t}\right)\leq$ $2\left(\frac{|E_{t}|}{|Q|}\right)^{\frac{1}{2^{n+1}[w]_{A_{\infty}}}}w(Q)$
$\leq$ $2\left(2e^{-c_{n}t}\right)^{\frac{1}{2^{n+1}[w]_{A_{\infty}}}}w(Q)$
$\leq$ $2^{1+\frac{1}{2^{n+1}}}e^{\frac{-c_{n}t}{2^{n+1}[w]_{A_{\infty}}}}w(Q).$
This concludes the proof of (1.5). To prove (1.6), it is enough to observe that
$\left\{x\in Q:Mf(x)-\operatorname*{ess\,inf}_{Q}Mf>\lambda,M^{\#}f(x)\leq\gamma\lambda\right\}\subseteq E_{\frac{1}{\gamma}},$
and apply (1.5). ∎
Proof of Theorem $1.$
Let $1\leq p<\infty$, $w\in A_{\infty}$ and let $Q$ be a cube. For each $t>0$, define
$E_{t}=\left\{x\in Q:\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}>t\right\}.$
By the layer-cake formula and (1.5), we have
$\left(\frac{1}{w(Q)}\int_{Q}\left(\frac{Mf(x)-\operatorname*{ess\,inf}_{Q}Mf}{M^{\#}f(x)}\right)^{p}w(x)\,dx\right)^{\frac{1}{p}}=$ $\left(p\int_{0}^{\infty}t^{p-1}\frac{w(E_{t})}{w(Q)}\,dt\right)^{\frac{1}{p}}$
$\leq$ $\left(c_{1}p\int_{0}^{\infty}t^{p-1}e^{-\frac{c_{2}t}{[w]_{A_{\infty}}}}\,dt\right)^{\frac{1}{p}}$
$=$ $c_{1}^{\frac{1}{p}}\frac{[w]_{A_{\infty}}}{c_{2}}\left(p\int_{0}^{\infty}s^{p-1}e^{-s}\,ds\right)^{\frac{1}{p}}$
$=$ $c_{1}^{\frac{1}{p}}\frac{[w]_{A_{\infty}}}{c_{2}}\left(p\Gamma(p)\right)^{\frac{1}{p}}$
$\leq$ $\frac{c_{1}}{c_{2}}[w]_{A_{\infty}}\Gamma(p+1)^{\frac{1}{p}},$
and we obtain the desired estimate by noting that $\Gamma(p+1)^{\frac{1}{p}}\lesssim p$ (see for instance [Grafakos, Corollary 6.2.5]). ∎
Remark 3.1.
Using similar arguments, we could also prove variants of Theorem 1.1 where the $L^{p}$ norm is replaced by Lorentz norms $L^{p,q}$. We do not pursue that direction.
The remainder of this section is devoted to the proof of Theorem 1.5. Within the proof, we will address some technical difficulties caused by the maximal operator, and then follow the lines of the proof of Theorem 1.2 in [OPRR].
Proof of Theorem 1.5.
In (1.4) (with $p=1$) we have already proved that
$\|Mf\|_{\operatorname{BLO}_{w}}\leq c_{n}[w]_{A_{\infty}}\|f\|_{\operatorname{BMO}},$
which provides the second inequality in (1.7). To prove the first inequality, we define
$X=\sup_{\|f\|_{\operatorname{BMO}}=1}\|Mf\|_{\operatorname{BLO}_{w}},$
where the supremum is taken over all functions $f$ such that $\|f\|_{\operatorname{BMO}}=1$ and MfMf is not identically infinite. We have to prove $w]_{A_{\infty}}\lesssim_{n}X$. Without loss of generality, we can assume that $X<\infty$.
Let $Q$ be a cube. We define the auxiliary weight
$v:=\left[M\left(\frac{w\chi_{Q}}{w_{Q}}\right)\right]^{\frac{1}{2}}.$
It is well known that $v\in A_{1}$ with $v]_{A_{1}}\leq c_{n}$ (see for instance [GCRdF]). Let us define $b:=\log^{+}v=\max(0,\log v)$, which satisfies $b\in\operatorname{BMO}$ and $\|b\|_{\operatorname{BMO}}\leq\tfrac{3}{2}\|\log v\|_{\operatorname{BMO}}\leq\rho_{n}$ (see [Grafakos, Proposition 6.1.4]). Using the definition of $X$, we obtain
$\|Mb\|_{\operatorname{BLO}_{w}}\leq\|b\|_{\operatorname{BMO}}X\leq\rho_{n}X,$
and hence,
$\frac{1}{w(Q)}\int_{Q}\left(Mb(x)-\operatorname*{ess\,inf}_{Q}Mb\right)w(x)\,dx\leq\rho_{n}X.$ (3.1)
We now estimate the essential infimum term. Fix $x\in Q$, and let $P$ be any cube with $x\in P$. By Jensen’s inequality,
$\frac{1}{|P|}\int_{P}|b(y)|\,dy=$ $\frac{1}{|P|}\int_{P}\log^{+}v(y)\,dy$
$\leq$ $\frac{1}{|P|}\int_{P}\log(v(y)+1)\,dy$
$\leq$ $\log\left(\frac{1}{|P|}\int_{P}v(y)\,dy+1\right)$
$\leq$ $\log\left(Mv(x)+1\right).$
Consequently, we have $Mb(x)\leq\log\left(Mv(x)+1\right)$ for every $x\in Q$. Using the $A_{1}$ condition and that $v\geq 1$ in $Q$, we have
$Mb(x)\leq\log([v]_{A_{1}}v(x)+1)\leq\log([v]_{A_{1}}+1)+\log v(x)\leq\log c_{n}+\log v(x)$
for almost every $x\in Q$. Therefore,
$\operatorname*{ess\,inf}_{z\in Q}Mb(z)\leq\log c_{n}+\operatorname*{ess\,inf}_{z\in Q}\log(v(z)).$
Using the Kolmogorov inequality together with the weak $(1,1)$ boundedness of the maximal operator (see [GrafakosClassic, p. 100]), we obtain
$\operatorname*{ess\,inf}_{z\in Q}Mb(z)\leq$ $\log c_{n}+\operatorname*{ess\,inf}_{z\in Q}\log(v(z))$
$\leq$ $\log c_{n}+\frac{1}{|Q|}\int_{Q}\log v$
$\leq$ $\log c_{n}+\log\left(\frac{1}{|Q|}\int_{Q}v\right)$
$=$ $\log c_{n}+\log\left(\frac{1}{|Q|}\int_{Q}M\left(\frac{w\chi_{Q}}{w_{Q}}\right)^{\frac{1}{2}}\right)$
$\leq$ $\log c_{n}+\log\left(2\left\|M\left(\frac{w\chi_{Q}}{w_{Q}}\right)\right\|_{L^{1,\infty}({\mathbb{R}}^{n},\frac{1}{|Q|})}^{\frac{1}{2}}\right)$
$\leq$ $\log c_{n}+\log c_{n}^{\prime}$
$=$ $\log\alpha_{n}.$
Note that
$\frac{1}{|Q|}\int_{Q}\log^{+}\left(\frac{w(x)}{\alpha_{n}^{2}w_{Q}}\right)w(x)\,dx=\frac{1}{|Q|}\int_{L_{Q}}\log^{+}\left(\frac{w(x)}{\alpha_{n}^{2}w_{Q}}\right)w(x)\,dx,$
where $L_{Q}=\{x\in Q:w(x)\geq\alpha_{n}^{2}w_{Q}\}$. Hence, for a.e. $x\in L_{Q}$ we have
$Mb(x)-\operatorname*{ess\,inf}_{z\in Q}Mb(z)\geq$ $Mb(x)-\log\alpha_{n}$
$\geq$ $b(x)-\log\alpha_{n}$
$=$ $\log v(x)-\log\alpha_{n}$
$=$ $\frac{1}{2}\log\left(\frac{1}{\alpha_{n}^{2}}M\left(\frac{w\chi_{Q}}{w_{Q}}\right)(x)\right)$
$\geq$ $\frac{1}{2}\log\left(\frac{1}{\alpha_{n}^{2}}\frac{w(x)}{w_{Q}}\right)$
$=$ $\frac{1}{2}\log^{+}\left(\frac{w(x)}{\alpha_{n}^{2}w_{Q}}\right).$
Combining this with (3.1), we obtain
$\frac{1}{2}\frac{1}{|Q|}\int_{Q}\log^{+}\left(\frac{w(x)}{\alpha_{n}^{2}w_{Q}}\right)w(x)\,dx=$ $\frac{1}{|Q|}\int_{L_{Q}}\frac{1}{2}\log^{+}\left(\frac{w(x)}{\alpha_{n}^{2}w_{Q}}\right)w(x)\,dx$
$\leq$ $\frac{1}{|Q|}\int_{L_{Q}}\left(Mb(x)-\operatorname*{ess\,inf}_{z\in Q}Mb(z)\right)w(x)\,dx$
$\leq$ $\frac{1}{|Q|}\int_{Q}\left(Mb(x)-\operatorname*{ess\,inf}_{z\in Q}Mb(z)\right)w(x)\,dx$
$\leq$ $\frac{w(Q)}{|Q|}\rho_{n}X.$
We conclude that
$\frac{1}{|Q|}\int_{Q}M(w\chi_{Q})(x)\,dx\simeq_{n}$ $\frac{1}{|Q|}\int_{Q}\left(1+\log^{+}\left(\frac{w(x)}{w_{Q}}\right)\right)w(x)\,dx$
$=$ $\frac{w(Q)}{|Q|}+\frac{1}{|Q|}\int_{Q}\log^{+}\left(\alpha_{n}^{2}\frac{w(x)}{\alpha_{n}^{2}w_{Q}}\right)w(x)\,dx$
$\leq$ $3\log\alpha_{n}\frac{w(Q)}{|Q|}+\frac{1}{|Q|}\int_{Q}\log^{+}\left(\frac{w(x)}{\alpha_{n}^{2}w_{Q}}\right)w(x)\,dx$
$\leq$ $3\log\alpha_{n}\frac{w(Q)}{|Q|}+2\frac{w(Q)}{|Q|}\rho_{n}X$
$\leq$ $C_{n}\frac{w(Q)}{|Q|}X,$
where we refer to [HP, Lemma 6.2] and [Stein] for the first equivalence, and in the last inequality we have used the bound $X\geq\frac{1}{2}$, which remains to be proved.
We now establish the inequality $X\geq\frac{1}{2}$. Let $H\subset{\mathbb{R}}^{n}$ be a half-space and let $Q$ be any cube intersecting both $H$ and its complement $H^{c}$. We may assume that
$w(Q\cap H)\geq\frac{1}{2}w(Q).$
Let $f=2\chi_{H}$. We have $\|\chi_{H}\|_{\operatorname{BMO}}=\frac{1}{2}$, and then $\|f\|_{\operatorname{BMO}}=1$. For $x\in H$, we have $M\chi_{H}(x)=1$ (since small cubes centered at $x$ are contained in $H$), and therefore $Mf(x)=2$. For $x\in H^{c}$, every cube containing $x$ intersects $H$ in at most half of its measure, and as the cube grows, the ratio tends to $\frac{1}{2}$. Consequently,
$M\chi_{H}(x)=\frac{1}{2}$
for each $x\in H^{c}$, and thus $Mf(x)=1$ on $H^{c}$. In particular,
$\operatorname*{ess\,inf}_{z\in Q}Mf(z)=1.$
We can now estimate the $\operatorname{BLO}_{w}$ seminorm of MfMf:
$Mf(x)-\operatorname*{ess\,inf}_{z\in Q}Mf(z)=\begin{cases}1&\text{if }x\in Q\cap H,\\ 0&\text{if }x\in Q\cap H^{c}.\end{cases}$
Hence,
$\frac{1}{w(Q)}\int_{Q}\left(Mf(x)-\operatorname*{ess\,inf}_{z\in Q}Mf(z)\right)w(x)\,dx=\frac{w(Q\cap H)}{w(Q)}\geq\frac{1}{2}.$
This proves that $X\geq\frac{1}{2}$, and therefore the proof is complete. ∎
We will now apply this characterization of the $A_{\infty}$ class to prove Corollary 1.6, which follows as a direct consequence of the quantitative inequalities established above.
Proof of Corollary 1.6.
The first part follows directly from Theorem $1.$ Let us now prove the second part. Let $w$ be a weight and assume that inequality (1.8) holds for every nonconstant locally integrable function $f$ such that MfMf is not identically infinite and $\|f\|_{\operatorname{BMO}}=1$. In particular, if $f\in\operatorname{BMO}$, then taking $p=1$ in (1.8) we have
$\|Mf\|_{\operatorname{BLO}_{w}}\leq C\|f\|_{\operatorname{BMO}}.$
By Theorem 1.5, this yields $w\in A_{\infty}$. In addition, by taking the supremum over $f$ we obtain,
$[w]_{A_{\infty}}\simeq_{n}$ $\sup_{f}\sup_{Q}\frac{1}{w(Q)}\int_{Q}\left(Mf(x)-\operatorname*{ess\,inf}_{Q}Mf\right)w(x)\,dx\leq C.$
This completes the proof. ∎
Acknowledgment
Part of this work was carried out while the author was visiting the Department of Mathematics at the Universidad de Buenos Aires. The author wishes to express his sincere gratitude to Ezequiel Rela for his hospitality during this visit.
Conflicts of Interest
The author has no conflicts of interest to declare.
Data Availability Statement
Data sharing is not applicable to this article, as no datasets were generated or analyzed during the current study.